# Signature matrices of membranes

### v1

@@ 5 Conclusions and future work
Furthermore, we propose novel closed formulas for signatures of piecewise bilinear and polynomial membranes. These membranes can be used for interpolating arbitrary image data, a necessary step in the iterated-integrals signature method. This should be useful for [48] which, instead of any interpolation, approximates two-parameter integrals via certain two-parameter sums.
Inspired by linear complexity of the two-parameter sums signature on diagonal matrix compositions [18, Theorem 4.5], we would like to improve our computational complexity for interpolated image data, Theorem 3.19.
@@
Is there an algorithm with linear time complexity in $m$ and $n$ to compute a level $k$ signature entry of a piecewise bilinear membranes of order $(m,n)$?
Already in the path framework, a signature matrix is not enough to recover the path, see [2, Example 6.8]. In particular, we can not recover a membrane from its signature matrix in general. An important property of the ‘right’ notion of signature is the ability of recovering membranes from all of its signature tensors, as it is done in [8] for paths. It is an open question whether our signature tensors have this property.
@@
Can we always recover membranes from their signature tensors?
One can also ask the following more special question:
Question 5.4.
@@
Again, this is true for paths, see [41].
We also remark that recovery with the two-parameter sums signature is always possible. For this compare [18, Theorem 2.20] and its proof.
Finally, we would like to understand the missing dimensions for membranes of very low order. Computations as shown in Figure 5 inspire the following conjecture:
Conjecture 5.5.
For $m,n\geq 2$, $\mathcal{P}_{d,m,n}$ has full dimension whenever $m+n>d$.
Note that these cases are not covered by Remark 4.8.
Question 5.6.
What are the dimensions of $\mathcal{P}_{d,n,m}$ for $m+n\leq d<mn$?
Acknowledgments
We thank Joscha Diehl for detailed discussions on the various notions of two-parameter signatures, Sven Groß for introducing us to interpolation literature, Ben Hollering for providing Example 4.14, Rosa Preiß for pointing out [14, Example 4.2], and Bernd Sturmfels for his help with Example 4.12. The authors acknowledge seed support for DFG CRC/TRR 388 “Rough Analysis, Stochastic Dynamics and Related Topics”.

### v2

@@ 5 Outlook and future work
We provide computational evidence for $d=k=3$ in Figure 7.
$m\!\setminus\!n$ 1 2 3 4 5
1 3 6 9 12 14
2 6 12 18 24 27
3 9 18 27
4 12 24
5 14 27
Figure 7: Dimensions of the Zariski closure of third level signature tensors of piecewise bilinear membranes of order $(m,n)$ in $\mathbb{R}^{3}$, where we omit the values once the full dimension has been reached.
Already in the path framework, a signature matrix is not enough to recover the path, see [bib:AFS2019, Example 6.8]. In particular, we can not recover a membrane from its signature matrix in general. An important property of the ‘right’ notion of signature is the ability of recovering membranes from all of its signature tensors, as it is done in [Chen1954IteratedIA] for paths. It is an open question whether our signature tensors have this property.
@@
Can we always recover membranes from its signature tensors?
One can also ask the following more special question:
@@
Again, this is true for paths, see [bib:PSS2019].
We also remark that recovery with the two-parameter sums signature is always possible. For this compare [DS22, Theorem 2.20] and its proof.
Acknowledgements
We thank Joscha Diehl for detailed discussions on various notions of two-parameter signatures, Sven Groß for introducing us to the interpolation literature, Rosa Preiß for pointing out [diehl2024signature, Example 4.2], and Ben Hollering and Bernd Sturmfels for their help with Examples 4.15 and 4.12.
Funding sources
The authors acknowledge support from DFG CRC/TRR 388 “Rough Analysis, Stochastic Dynamics and Related Fields”, Project A04.
[…]